# Execution Guided Line-by-Line Code Generation

### v1

@@ Abstract
We present a novel approach to neural code generation that incorporates real-time execution signals into the language model generation process. While large language models (LLMs) have demonstrated impressive code generation capabilities, they typically do not utilize execution feedback during inference, a critical signal that human programmers regularly leverage. Our method, Execution-Guided Classifier-Free Guidance (EG-CFG), dynamically incorporates execution signals as the model generates code, providing line-by-line feedback that guides the generation process toward executable solutions. EG-CFG employs a multi-stage process: first, we conduct beam search to sample candidate program completions for each line; second, we extract execution signals by executing these candidates against test cases; and finally, we incorporate these signals into the prompt during generation. By maintaining consistent signals across tokens within the same line and refreshing signals at line boundaries, our approach provides coherent guidance while preserving syntactic structure. Moreover, the method naturally supports native parallelism at the task level in which multiple agents operate in parallel, exploring diverse reasoning paths and collectively generating a broad set of candidate solutions. Our experiments across diverse coding tasks demonstrate that EG-CFG significantly improves code generation performance compared to standard approaches, achieving state-of-the-art results across various levels of complexity, from foundational problems to challenging competitive programming tasks. Our code is available at: https://github.com/boazlavon/eg_cfg
@@ 1 Introduction
Large language models (LLMs) have recently demonstrated remarkable code generation capabilities, significantly advancing performance in tasks such as general programming problems chen2021evaluating ; jiang2024survey ; austin2021program , competitive coding challenges li2022competition , and real-world software engineering tasks hou2024large . Despite this progress, current LLM-based code generation methods primarily rely on pattern recognition derived from static representations of code rather than explicitly modeling code execution at runtime chen2021evaluating ; austin2021program . Consequently, the generated programs often appear to be correct superficially, but fail to execute correctly on actual inputs, reflecting a critical gap between learned syntactic patterns and genuine executability.
@@
Figure 1: MBPP and MBPP-ET performance. Our method, EG-CFG, achieves a new SOTA using an open-source LLM, surpassing methods utilizing closed-source ones.
Figure 2: HumanEval and HumanEval-ET performance. Our method, EG-CFG, achieves a new SOTA on HumanEval-ET benchmark.
As depicted in Figure 2, our Execution-Guided Classifier-Free Guidance (EG-CFG) approach achieves state-of-the-art performance on the MBPP and MBPP-ET benchmarks austin2021program , significantly outperforming existing methods. Using the open-source DeepSeek-V3-0324 model liu2024deepseek , EG-CFG attains 96.6% accuracy on MBPP and 73.0% on MBPP-ET, surpassing previous leading approaches such as QualityFlow hu2025qualityflow (94.2%), MetaGPT metagpt (87.7%), and LPW lei2024planning (84.8% on MBPP, 65.8% on MBPP-ET), all of which utilized leading closed-source models. On the HumanEval benchmark (Figure 2), EG-CFG achieves highly competitive results compared to existing state-of-the-art methods and establishes a new state-of-the-art on the HumanEval-ET benchmark, reaching 87.19% accuracy and surpassing LPW’s 84.8% accuracy achieved with GPT-4o. Lastly, EG-CFG achieves a new state-of-the-art on the CodeContests benchmark (Figure 3) with 58.18% accuracy using DeepSeek-V3-0324, demonstrating its effectiveness on challenging competitive programming problems.
Figure 3: CodeContests performance. Our method, EG-CFG, achieves a new SOTA.
The superior performance of EG-CFG on MBPP-ET and HumanEval-ET highlights that our method not only generates accurate code but also significantly enhances robustness and reliability under complex and extended test scenarios.
Summarizing, our main contributions are: (i) a framework for dynamically generating code while executing fragments of the code and using the execution traces to guide the generation process, (ii) introducing native parallelism at the task level that is not achievable in the sequential iterative refinement methods, (iii) using CFG in order to generate code that is conditioned on the execution feedback, and (iv) obtaining new state of the art results on the MBPP, MBPP-ET, HumanEval-ET, and CodeContests benchmarks using open-source models, outperforming previous methods that are based on the leading closed-source models.
@@ 2 Related Work
Program synthesis has long served both as a means to evaluate LLMs’ capabilities chen2021evaluating ; hendrycks2measuring and as a key goal of the research community to automate code generation waldinger1969prow ; manna1971toward . Modern LLMs are typically assessed via automated benchmarks: given a coding problem description, the model generates code that is then validated on a set of test cases (unit-test). In this work, we focus on three code generation benchmarks: MBPP austin2021program , HumanEval chen2021evaluating , and CodeContests li2022competition , along with the enhanced variants MBPP-ET and HumanEval-ET dong2025codescore . MBPP combines crowd-sourced tasks and math-based problems (amini2019mathqa, ), HumanEval includes hand-written Python challenges, and CodeContests features competitive programming problems requiring advanced algorithmic reasoning.
@@
While some approaches use a single LLM to solve coding problems, agentic methods employ multiple LLM instances to tackle a single task. Works such as MetaGPT metagpt , introduce a framework in which a given task is split into multiple procedures, each assigned to a dedicated LLM agent. AgentCoder huang2023agentcoder extends these multi-agent frameworks by employing both a test-designer agent and an executor agent, effectively integrating the Self-Debugging paradigm chen2023teaching into an agentic workflow and demonstrating the benefits of execution feedback. Current agentic workflows, such as MapCoder islam2024mapcoder and LPW lei2024planning , typically rely on discrete cycles of sequential agent interactions or distinct refinement phases, inherently limiting their ability to fully leverage the power of parallelism within a single task. In contrast, EG-CFG breaks this barrier by introducing native parallelism at the task level. Multiple agents run concurrently on the same task, each with a different configuration, simultaneously exploring diverse reasoning paths without the constraints of sequential cycles, fully leveraging the power of parallelism.
Although code generation methods have advanced dramatically, they all rely on naive decoding methods such as temperature or top-p sampling holtzman2019curious . In particular, execution-feedback techniques generate a complete solution, execute it, and then refine the code. For example, the model is prompted to produce a function, its implementation is tested against unit tests, and the resulting feedback is appended to the original prompt. This process repeats iteratively until a correct solution emerges. This approach contrasts with recent advances in LLM inference techniques, especially guidance methods such as Classifier Guidance dhariwal2021diffusion and Classifier-Free Guidance (CFG) ho2022classifier . These methods condition generation on external constraints or classifier signals, directing the model to sample tokens from multiple contrastive distributions. Although CFG has shown strong performance, its guidance signals remain static, predefined and fixed throughout the sampling process sanchez2024stay .
As far as we can ascertain, guidance methods such as CFG have yet to be applied at scale with dynamic, execution-driven reasoning. This work bridges this gap by demonstrating that a single open-source model, augmented with execution feedback and a novel token-sampling strategy, outperforms the state of the art across multiple widely adopted coding benchmarks at different complexity levels - from foundational problems to challenging competitive programming tasks. Our contribution differs from prior work in multiple key aspects, including:
@@
    Prior approaches generate entire code blocks (or sub-blocks) and using the execution feedback of the entire block; by contrast, our approach gradually generates the solution line-by-line by sampling and executing multiple candidate continuations programs at each step.
@@ 3 Method
We present Execution-Guided Context-Free Guidance (EG-CFG), a novel inference-time decoding method for neural code generation that explicitly integrates dynamic execution signals into the autoregressive generation process.
A programming task is represented by the templated prompt composed of four components:
$p_{\text{inst}}=(p_{0},p_{\text{task}},T,f_{\text{name}}),$ (1)
where $p_{0}$ is the natural-language instruction template, $p_{\text{task}}$ is a textual definition of the programming task, $T=\{t_{j}\}_{j=1}^{|T|}$ is the set of test cases, and $f_{\text{name}}$ is the target Python function name. The goal is to generate executable Python code $w^{*}=[w_{0}^{*},w_{1}^{*},\dots,w_{N-1}^{*}]$ that solves the task correctly, formally satisfying:
@@
At each step $i$, given a previously generated token sequence $w_{<i}$, the LLM $M$ assigns a probability distribution $M(w_{i}\mid p_{\text{inst},w_{<i}})$, which is conditioned on the instruction prompt $p_{\text{inst}}$. This generation proceeds iteratively until reaching a maximum token length $N_{\text{max}}$ or an end-of-code token.
@@
We first run the model to produce an initial solution $p_{\text{pre}}$ based on $p_{\text{inst}}$
@@
until we identify the beginning of the final executable solution block within, by locating the occurrence of a special start-of-code token, which is defined in $p_{0}$. This position is marked as $i_{\text{solution}}$. We note that $p_{\text{pre}}$ contains reasoning tokens, and other tokens that precede the final start-of-code token. Another index that we define $p_{\text{pre}}$ is the index $i_{\text{dyn}}$ of the last token of the last few example inside the template $p_{\text{inst}}$. We used this location to inject future signals into the prompt.
The prompt $p_{\text{sol}}$ that we pass to the LLM is created autoregressively by aggregating executable code tokens into $p_{\text{pre}}$. This prompt contains partial solutions that are updated until the solution is formed.
@@ 3.1 Dynamic Execution Feedback
Our dynamic execution feedback explicitly generates multiple candidate continuations based on a partially completed solution. Specifically, given $p_{\text{sol}}$ we generate a set of candidate continuations using beam-search decoding. Each candidate explicitly extends the current solution by $d$ additional lines of code, capturing meaningful variations in potential solutions and providing a granular basis for execution-based guidance. Formally, given parameters specifying the number of candidates $s$, completion horizon $d$, and sampling temperature $t$, the beam search sampling is performed to obtain $s$ candidates:
$c^{j}\sim M(p_{\text{sol}};d,t),\quad j=1,\dots,s$ (4)
where each candidate $c^{j}$ explicitly represents a plausible continuation of the next $d$ lines of code. These candidates form the basis for generating detailed execution signals used to guide subsequent inference steps.
@@
Since generated candidate continuations may contain incomplete or syntactically invalid Python code, we explicitly extract their executable components using syntactic validation via Abstract Syntax Tree (AST) parsing. Formally, for each candidate $c^{j}$, we apply the executable extraction function:
@@
This extraction function iteratively attempts to parse each candidate $c^{j}$ as follows: (1) attempt parsing $c^{j}$ (2) if unsuccessful, append a Python pass statement to the last line and retry; (3) if still unsuccessful, iteratively remove the last line and retry. This ensures minimal modifications for syntactic validity.
After extraction, we apply uniqueness filtering to remove duplicates:
@@
Our approach is agnostic to the precise structure of the execution feedback $e^{j,m}$. In our implementation, execution feedback specifically takes the form of execution traces: a structured representation of a program’s runtime behavior, capturing detailed step-by-step information during execution. Specifically, an execution trace includes executed lines of code, variable states (values and types), function invocations, return values, and detailed descriptions of runtime errors, if they occur. This structured log provides comprehensive insight into the correctness and behavior of executed code, serving as a precise basis for dynamic feedback in our inference framework.
@@
where each tuple consists of a candidate completion $\hat{c}^{j}$, a test case $t_{m}$, and its corresponding execution trace $e^{j,m}$.
Now we form a new prompt naming this prompt $p_{\text{dyn}}$ as dynamic signal prompt:
@@
An example of the obtained prompt is shown in Figure 4.
@@ 3.2 Classifier-Free Guidance (CFG)
Inspired by (ho2022classifier, ), we utilize CFG to explicitly guide token generation by interpolating between two probability distributions: (i) an unconditional (prior) distribution based on the evolving solution prompt $p_{\text{sol}}$, and (ii) a conditional distribution based on dynamic signal prompt $p_{\text{dyn}}$ which incorporates execution feedback. Formally, for each token $w_{i}$, the CFG distribution is computed as:

### v2

# spacing=nonfrench

@@ Abstract
We present a novel approach to neural code generation that incorporates real-time execution signals into the language model generation process. While large language models (LLMs) have demonstrated impressive code generation capabilities, they typically do not utilize execution feedback during inference, a critical signal that human programmers regularly leverage. Our method, Execution-Guided Classifier-Free Guidance (EG-CFG), dynamically incorporates execution signals as the model generates code, providing line-by-line feedback that guides the generation process toward executable solutions. EG-CFG employs a multi-stage process: first, we conduct beam search to sample candidate program completions for each line; second, we extract execution signals by executing these candidates against test cases; and finally, we incorporate these signals into the prompt during generation. By maintaining consistent signals across tokens within the same line and refreshing signals at line boundaries, our approach provides coherent guidance while preserving syntactic structure. Moreover, the method naturally supports native parallelism at the task level in which multiple agents operate in parallel, exploring diverse reasoning paths and collectively generating a broad set of candidate solutions. Our experiments across diverse coding tasks demonstrate that EG-CFG significantly improves code generation performance compared to standard approaches, achieving state-of-the-art results across various levels of complexity, from foundational problems to challenging competitive programming and data science tasks. Our code is available at: https://github.com/boazlavon/eg_cfg
@@ 1 Introduction
Large language models (LLMs) have recently demonstrated remarkable code generation capabilities, significantly advancing performance in tasks such as general programming problems chen2021evaluating ; jiang2024survey ; austin2021program , competitive coding challenges li2022competition , and real-world software engineering tasks hou2024large . However, current LLM-based code generation methods primarily rely on pattern recognition derived from static representations of code rather than explicitly modeling code execution at runtime chen2021evaluating ; austin2021program . Consequently, the generated programs often appear to be correct superficially, but fail to execute correctly on actual inputs, reflecting a critical gap between learned syntactic patterns and genuine executability.
@@
Figure 1: MBPP & MBPP-ET performance. EG-CFG (DeepSeek-V3) sets a new state-of-the-art results.
Figure 2: HumanEval & HumanEval-ET performance. EG-CFG (DeepSeek-V3) matches the state-of-the-art on HumanEval and sets a new state-of-the-art on HumanEval-ET.
Figure 3: DS-1000 performance. EG-CFG (DeepSeek-V3) achieves new state-of-the-art, surpassing GPT-4.
Figure 4: CodeContests performance. EG-CFG (DeepSeek-V3) sets a new state-of-the-art, outperforming GPT-4 and GPT-4o methods.
As depicted in Figure 1, our Execution-Guided Classifier-Free Guidance (EG-CFG) approach achieves state-of-the-art performance on the MBPP and MBPP-ET benchmarks austin2021program , significantly outperforming existing methods. Using the open-source DeepSeek-V3-0324 model liu2024deepseek , EG-CFG attains 96.6% accuracy on MBPP and 73.0% on MBPP-ET, surpassing previous leading approaches such as QualityFlow hu2025qualityflow (94.2%), MetaGPT metagpt (87.7%), and LPW lei2024planning (84.8% on MBPP, 65.8% on MBPP-ET), all of which utilized leading closed-source models. On the HumanEval benchmark (Figure 2), EG-CFG achieves state-of-the-art accuracy of 99.4%, matching LDB zhong2024debug and establishes a new state-of-the-art on the HumanEval-ET benchmark, reaching 89.02% accuracy and surpassing LPW’s 84.8% accuracy achieved with GPT-4o. Furthermore, EG-CFG establishes a new state-of-the-art on DS-1000 (Figure 3), a domain-specific benchmark focused on challenging data science problems, achieving 69.9% accuracy. Lastly, EG-CFG achieves a new state-of-the-art on the CodeContests benchmark (Figure 4) with 60.6% accuracy using DeepSeek-V3-0324, demonstrating its effectiveness on challenging competitive programming problems.
The superior performance of EG-CFG on MBPP-ET and HumanEval-ET highlights that our method not only generates accurate code but also significantly enhances robustness and reliability under complex and extended test scenarios. Summarizing, our main contributions are: (i) a framework for dynamically generating code while executing fragments of the code and using the execution traces to guide the generation process, (ii) introducing native parallelism at the task level that is not achievable in the sequential iterative refinement methods, (iii) using CFG in order to generate code that is conditioned on the execution feedback, and (iv) obtaining new state-of-the-art results on the MBPP, MBPP-ET, HumanEval-ET, DS-1000 and CodeContests benchmarks using open-source models, outperforming previous methods that are based on the leading closed-source models.
@@ 2 Related Work
Program synthesis has long served both as a means to evaluate LLMs’ capabilities chen2021evaluating ; hendrycks2020measuring and as a key goal of the research community to automate code generation waldinger1969prow ; manna1971toward . Modern LLMs are typically assessed via automated benchmarks: given a coding problem description, the model generates code that is then validated on a set of test cases (unit-test). In this work, we focus on six code generation benchmarks: MBPP austin2021program , HumanEval chen2021evaluating , DS-1000 Lai2022DS1000 and CodeContests li2022competition , along with the extended variants MBPP-ET and HumanEval-ET dong2025codescore . MBPP combines crowd-sourced tasks and math-based problems (amini2019mathqa, ), HumanEval includes hand-written Python challenges, DS-1000 presents challenging data science problems involving libraries like Pandas and NumPy, and CodeContests features competitive programming problems requiring advanced algorithmic reasoning.
@@
While some approaches use a single LLM to solve coding problems, agentic methods employ multiple LLM instances to tackle a single task. Works such as MetaGPT metagpt introduce a framework in which a given task is split into multiple procedures, each assigned to a dedicated LLM agent. AgentCoder huang2023agentcoder extends these multi-agent frameworks by employing both a test-designer agent and an executor agent, effectively integrating the Self-Debugging paradigm chen2023teaching into an agentic workflow and demonstrating the benefits of execution feedback. Current agentic workflows, such as MapCoder islam2024mapcoder and LPW lei2024planning , typically rely on discrete cycles of sequential agent interactions or distinct refinement phases, inherently limiting their ability to fully leverage the power of parallelism within a single task. In contrast, EG-CFG breaks this barrier by introducing native parallelism at the task level. Multiple agents run concurrently on the same task, each with a different configuration, simultaneously exploring diverse reasoning paths without the constraints of sequential cycles, fully leveraging the power of parallelism.
Although code generation methods have advanced dramatically, they all rely on naive decoding methods such as temperature or top-p sampling holtzman2019curious . In particular, execution-feedback techniques generate a complete solution, execute it, and then refine the code. For example, the model is prompted to produce a function, its implementation is tested against unit tests, and the resulting feedback is appended to the original prompt. This process repeats iteratively until a correct solution emerges. This approach contrasts with recent advances in LLM inference techniques, especially guidance methods such as Classifier Guidance dhariwal2021diffusion and Classifier-Free Guidance (CFG) ho2022classifier ; sanchez2023stay . These methods condition generation on external constraints or classifier signals, directing the model to sample tokens from multiple contrastive distributions. Although CFG has shown strong performance, its guidance signals remain static, predefined and fixed throughout the sampling process sanchez2024stay .
As far as we can ascertain, guidance methods such as CFG have yet to be applied at scale to dynamic, execution-driven reasoning. This work bridges this gap by demonstrating that a single open-source model, augmented with execution feedback and a novel token-sampling strategy, outperforms the state-of-the-art across multiple widely adopted coding benchmarks at different complexity levels - from foundational problems to challenging competitive programming and data sceince tasks. Our contribution differs from prior work in multiple key aspects, including:
@@
    Prior approaches generate entire code blocks (or sub-blocks) and use the execution feedback of the entire block; by contrast, our approach gradually generates the solution line-by-line by sampling and executing multiple candidate continuation programs at each step.
@@ 3 Method
We present Execution-Guided Classifier-Free Guidance (EG-CFG), a novel inference-time decoding method for neural code generation that explicitly integrates dynamic execution signals into the autoregressive generation process. A programming task $\tau$ is represented by three components:
$\tau=(p_{\text{task}},T,f_{\text{name}}),$ (1)
where $p_{\text{task}}$ is a textual definition of the programming task, $T=\{t_{j}\}_{j=1}^{|T|}$ is the set of test cases, and $f_{\text{name}}$ is the target Python function name. The goal is to generate executable Python code $w^{*}=[w_{0}^{*},w_{1}^{*},\dots,w_{N-1}^{*}]$ that solves the task correctly, formally satisfying:
@@
We build the instruction prompt by incorporating the task components into the prompt template:
$p_{\text{inst}}=\textsc{BuildInstructionPrompt}(p_{0},\tau)$ (3)
At each step $i$, given a previously generated token sequence $w_{<i}$, the LLM $M$ assigns a probability distribution $M(w_{i}\mid p_{\text{inst},w_{<i}})$, which is conditioned on the instruction prompt $p_{\text{inst}}$. This generation proceeds iteratively until reaching a maximum token length $N_{\text{max}}$ or an end-of-code token. We build the instruction prompt by injecting the instruction task to the prompt template. We first run the model to produce an initial output sequence $p_{\text{pre}}$ based on $p_{\text{inst}}$:
@@
The generated sequence $p_{\text{pre}}$ contains reasoning tokens that precede the final executable solution. We identify the beginning of this solution block at index $i_{\text{solution}}$ by locating a special start-of-code token, which is defined in $p_{0}$. We then define the prefix of all preceding tokens, $p_{\text{pre-sol}}$, as:
$p_{\text{pre-sol}}=p_{\text{pre}}[:i_{\text{solution}}]$ (5)
Additionally, we define an index $i_{\text{dyn}}$ that marks the end of the final few-shot example within the instruction prompt $p_{\text{inst}}$, which is used to inject future signals into the prompt.
The prompt $p_{\text{sol}}$ that we pass to the LLM is constructed autoregressively by aggregating executable code tokens into $p_{\text{pre-sol}}$ as formalized later in  Equation 14 of  section 3.3. This prompt contains partial solutions that are progressively updated until the solution is formed.
@@ 3.1 Dynamic Execution Feedback
Our dynamic execution feedback explicitly generates multiple candidate continuations based on a partially completed solution. Specifically, given $p_{\text{sol}}$ we generate a set of candidate continuations using beam-search decoding. Each candidate explicitly extends the current solution by $d$ additional lines of code, capturing meaningful variations in potential solutions and providing a granular basis for execution-based guidance. Formally, given parameters specifying the number of candidates $s$, completion horizon $d$, and sampling temperature $t$, the beam search sampling is performed to obtain $s$ candidates.
$w^{i}_{c_{j}}\sim M(w_{i}\mid p_{\text{sol}},w^{<i}_{c_{j}};t)\ \ \text{until }\ \mathrm{CountLines}(c_{j})\geq d,\ \ c_{j}=[w^{0}_{c_{j}},\cdots,w^{i}_{c_{j}}],\ \ j=1,\dots,s$ (6)
Each candidate $c^{j}$ is generated autoregressively with a stopping condition triggered after $d$ newline characters, representing a plausible continuation of the next $d$ lines of code. These candidates form the basis for generating detailed execution signals used to guide subsequent inference steps.
@@
To handle potentially invalid candidate continuations, we extract executable components via Abstract Syntax Tree (AST) parsing. Formally, for each candidate $c^{j}$, we apply the executable extraction function:
@@
This extraction function iteratively attempts to parse each candidate $c^{j}$ as follows: (1) attempt parsing $c^{j}$ (2) if unsuccessful, append a Python pass statement to the last line and retry; (3) if still unsuccessful, iteratively remove the last line and retry. This ensures minimal modifications for syntactic validity. After extraction, we apply uniqueness filtering to remove duplicates:
@@
Our approach is agnostic to the precise structure of the execution feedback $e^{j,m}$. In our implementation, execution feedback specifically takes the form of execution traces: a structured representation of a program’s runtime behavior, capturing detailed step-by-step information during execution.
Specifically, we use a custom debugger for structured execution traces. A trace $e^{j,m}$ is defined as a sequence of $N^{j,m}$ runtime events, $\varepsilon_{k}$, resulting from the execution of a program $\hat{c}^{j}$ with input $t_{m}$:
$e^{j,m}=\left[\varepsilon_{k}\right]_{k=1}^{N^{j,m}},\quad\varepsilon_{k}=\left(E_{k},\ell_{k},v_{k},\tau_{k},r_{k},x_{k}\right)$ (10)
Each event $\varepsilon_{k}$ is a tuple containing the event type $E_{k}\in\{\texttt{call},\texttt{line},\texttt{return},\texttt{exception}\}$, the source line number $\ell_{k}$, mappings from variable names to their values ($v_{k}$) and types ($\tau_{k}$), the return value $r_{k}$, and any exception details $x_{k}$. This structured representation provides comprehensive insight into the correctness and behavior of executed code, serving as a precise basis for dynamic feedback in our inference framework.
@@
where each tuple consists of a candidate completion $\hat{c}^{j}$, a test case $t_{m}$, and its corresponding execution trace $e^{j,m}$. Now we form a new prompt naming this prompt $p_{\text{dyn}}$ as dynamic signal prompt:
@@
An example of the obtained prompt is shown in appendix Appendix B.
@@ 3.2 Classifier-Free Guidance (CFG)
Inspired by (ho2022classifier, ; sanchez2023stay, ), we utilize CFG to explicitly guide token generation by interpolating between two probability distributions: (i) an unconditional (prior) distribution based on the evolving solution prompt $p_{\text{sol}}$, and (ii) a conditional distribution based on dynamic signal prompt $p_{\text{dyn}}$ which incorporates execution feedback. Formally, for each token $w_{i}$, the CFG distribution is computed as: